# RadioDiff-FS: Physics-Informed Manifold Alignment in Few-Shot Diffusion Models for High-Fidelity Radio Map Construction

### 19 Mar 2026 v1

@@ Abstract
Radio maps (RMs) provide spatially continuous propagation characterizations essential for 6G network planning, but high-fidelity RM construction remains challenging. Rigorous electromagnetic solvers incur prohibitive computational latency, while data-driven models demand massive labeled datasets and generalize poorly from simplified simulations to complex multipath environments. This paper proposes RadioDiff-FS, a few-shot diffusion framework that adapts a pre-trained main-path generator to multipath-rich target domains with only a small number of high-fidelity samples. The adaptation is grounded in a theoretical decomposition of the multipath RM into a dominant main-path component and a directionally sparse residual. This decomposition shows that the cross-domain shift corresponds to a bounded and geometrically structured feature translation rather than an arbitrary distribution change. A Direction-Consistency Loss (DCL) is then introduced to constrain diffusion score updates along physically plausible propagation directions, suppressing phase-inconsistent artifacts that arise in the low-data regime. Experiments show that RadioDiff-FS reduces NMSE by 59.5% on static RMs and by 74.0% on dynamic RMs relative to the vanilla diffusion baseline, achieving an SSIM of 0.9752 and a PSNR of 36.37 dB under severely limited supervision.
@@ I Introduction
Figure 1: Illustration of the few-shot in RM.
@@
    We establish a rigorous theoretical decomposition of the multipath RM into a dominant main-path component and a directionally sparse residual. This formulation provides a physical inductive bias demonstrating that the cross-domain transition is characterized by learnable geometric shifts rather than chaotic distribution changes, and lays the theoretical foundation for efficient few-shot adaptation.
@@
    We design a Direction-Consistency Loss that functions as a lightweight manifold projector within the diffusion process. By constraining score function updates to physically plausible propagation directions, DCL penalizes gradient components representing mode hallucinations, ensuring that synthesized maps maintain global spatial coherence while accurately recovering fine-grained diffraction and reflection details.
@@ II-A RM Construction Methods
Initial studies, such as RadioUNet [10], demonstrated that convolutional encoder–decoder architectures can produce accurate 2D pathloss maps from structured urban layouts, significantly reducing inference latency compared with ray tracing. Subsequent works like PMNet improved generalization and sample efficiency through advanced vision architectures. Although these discriminative models remain computationally efficient and interpretable, they typically require dense supervision and are difficult to adapt to out-of-distribution environments or scenarios with sparse sensor coverage. Moreover, most of these methods focus on planar pathloss estimation, ignoring vertical structure and multi-metric coupling [26, 22]. To overcome the limitations of pointwise regression under sparse measurements, conditional generative models have been explored. RME-GAN [35] and DeepREM [2] propose using conditional generative adversarial networks and completion autoencoders to reconstruct dense RMs from limited measurements. These models integrate global propagation patterns and local texture variations to enhance estimation performance. While they significantly outperform interpolation baselines, they still rely heavily on simulation data and do not incorporate electromagnetic consistency into the learning process. Graph neural networks (GNNs) offer another route by encoding structural priors derived from building layouts and sensor topology [3]. Recent methods formulate sparse measurements as nodes and use edge structures informed by connectivity or spatial correlation to facilitate map completion. RadioGAT [11] extends this by capturing spatial–spectral dependencies across multiple frequency bands. These models offer improved robustness to sampling irregularities and can generalize under semi-supervised settings. However, without explicit electromagnetic constraints, their behavior around complex propagation features such as diffraction edges and shadow zones remains difficult to guarantee.
@@
Diffusion models have recently emerged as a powerful alternative due to their generative structure and sample flexibility. RadioDiff [22] models RM generation as a conditional diffusion process and introduces attention mechanisms and fast Fourier modules to enhance feature extraction in dynamic environments. RadioDiff-3D builds upon this by incorporating 3D convolutional networks and utilizing the UrbanRadio3D dataset, which contains multi-height, multi-parameter RMs including pathloss, direction-of-arrival (DoA), and time-of-arrival (ToA) [23]. These models achieve state-of-the-art performance in high-dimensional map estimation. However, without embedding physical priors, they may fail to reconstruct physically meaningful structures in unseen domains, especially in low-data regimes. To address this, physics-informed models have been introduced to embed electromagnetic theory into neural estimators. PEFNet integrates the volume integral equation into the learning objective, constraining the relationship between incident and total electric fields. More advanced methods, such as PhyRMDM [8] and RadioDiff-$k^{2}$ [24], incorporate Helmholtz residuals, source terms, and boundary conditions into diffusion model training. RadioDiff-$k^{2}$ in particular establishes a theoretical correspondence between wave propagation singularities and negative wave-number domains, enabling the network to identify and reconstruct diffraction-sensitive structures via a two-stage generative process. These approaches significantly improve physical plausibility and generalization, but they often assume access to accurate simulations or well-calibrated measurements, which remain costly and difficult to obtain at scale. Task-specific strategies such as RM inpainting based on propagation priority and radio depth maps have also been proposed to address areas where sensing is physically restricted [20]. In parallel, the development of large-scale datasets such as UrbanRadio3D has greatly expanded the coverage of height, geometry, and multi-metric labeling, providing an essential benchmark for high-fidelity 3D×3D RM construction. Nevertheless, the cost and time required for data generation via ray tracing or controlled field measurements remain prohibitive in most deployment scenarios, and the resulting models often exhibit limited cross-region transferability.
@@ II-C Denoising Diffusion Model
where $f(\bm{x},t)$ denotes the drift term, $g(t)$ is a time-dependent diffusion coefficient, and $d\bm{w}$ represents a standard Wiener process. As $t$ increases, the perturbed distribution $p_{t}(\bm{x})$ asymptotically approaches an isotropic Gaussian. The generative process involves solving the corresponding reverse-time SDE:
@@
$s_{\theta}(\bm{x},t)=-\frac{\epsilon_{\theta}(\bm{x},t)}{\sqrt{1-\bar{\alpha}_{t}}},$ (6)
demonstrating DDPM as a discretized instance of score-based continuous models under a variance-preserving schedule.
@@ III System Model and Problem Formulation
We consider a wireless propagation scenario over a discretized planar region represented as an $N\times N$ uniform grid. Each cell in this grid corresponds to a spatial resolution element of width $\Delta$, and is assumed small enough such that the pathloss within it can be treated as constant. A RM over this region is thus defined as a matrix $\bm{P}\in\mathbb{R}^{N\times N}$, where each entry $P(i,j)$ denotes the pathloss at the center of cell $(i,j)$, measured at a fixed carrier frequency $f_{c}$ with corresponding wavelength $\lambda=c/f_{c}$, where $c$ denotes the speed of light in free space. A single transmitter, modeled as a point radiator, is positioned at $\bm{r}=\langle d_{x},d_{y},d_{z}\rangle$, with $(d_{x},d_{y})$ denoting its horizontal coordinates in the 2D region and $d_{z}>0$ indicating its height above the ground. The transmitter emits electromagnetic energy isotropically or according to a specified antenna gain pattern, and serves as the sole radiation source in the modeled scenario. While we focus on the single-transmitter setting, the formulation can be extended to multi-antenna or multi-site deployments by considering multiple independent transmitter locations. The surrounding environment is composed entirely of static structures, such as buildings, whose geometric layout is captured by a binary occupancy matrix $\bm{H}\in{0,1}^{N\times N}$. An entry $H(i,j)=1$ indicates that the cell at location $(i,j)$ is fully occupied by a non-penetrable obstacle, while $H(i,j)=0$ represents free space. These obstacles are assumed to fully block direct wave propagation. Given a transmitter position $r$ and static environment $\bm{H}$, the goal is to compute or estimate the pathloss field $\bm{P}_{\star}(\cdot,;\bm{r})=\mathcal{F}(\bm{H};\bm{r})$, where $\mathcal{F}$ denotes the environment-to-field operator. In this work, we consider two distinct types of RMs. The first, referred to as the static RM (SRM), models only the dominant propagation path and captures large-scale attenuation effects such as free-space spreading and shadowing due to occlusions. The second, referred to as the multi-path RM (MRM), provides a more detailed characterization of the propagation field by incorporating multipath effects, including reflections and diffractions induced by the surrounding static geometry. The MRM thus offers a more physically complete view of the underlying electromagnetic field, particularly in non-line-of-sight regions.
@@
To support the construction of these RMs, we assume the availability of two labeled datasets. The first, denoted as $\mathcal{D}_{\mathrm{SRM}}=\{\bm{H}_{k},\bm{r}_{k},\bm{P}_{k}^{\mathrm{SRM}}\}_{k=1}^{K}$, contains a large number of examples generated under the SRM setting. These samples are efficient to compute and cover a wide variety of layouts and transmitter placements. The second dataset, denoted as $\mathcal{D}_{\mathrm{MRM}}=\{\bm{H}_{\ell},\bm{r}_{\ell},\bm{P}_{\ell}^{\mathrm{MRM}}\}_{\ell=1}^{L}$, contains a small number of high-fidelity samples capturing the effects of multipath propagation. Due to the computational overhead associated with accurate multipath modeling, the size of this dataset is limited, i.e., $L\ll K$. The objective of the above problem can be formulated as follows.
@@
$\min_{\bm{\theta}}\sum_{\ell}^{L}\frac{1}{L}\mathcal{L}\left(\bm{P}_{\ell}^{\mathrm{MRM}},\hat{\bm{P}}_{\ell}^{\mathrm{MRM}}\right),$ (15)
$s.t.\hat{\bm{P}}_{\ell}^{\mathrm{MRM}}=\bm{\mu}_{\bm{\theta}}\left(\bm{H},\bm{r}|\mathcal{D}_{\mathrm{SRM}},\mathcal{D}_{\mathrm{MRM}}\right),$ (15a)
where $\bm{\theta}$ is trainable parameter of neural network $\bm{\mu}(\cdot)$.
@@ IV-A Main-Path Decomposition
Consider a bounded planar domain $\Omega\subset\mathbb{R}^{2}$ and a fixed transmitter–environment configuration. At a receiver location $\bm{x}\in\Omega$, let $\mathcal{P}(\bm{x})$ denote a finite set of propagation paths induced by the environment. Under the power-incoherent superposition model commonly adopted by RM construction, the received power field is expressed as the nonnegative sum
@@
where $P_{p}(\bm{x})$ aggregates geometric spreading and interaction-dependent losses along path $p$.
Let $p^{\star}(\bm{x})\in\arg\max_{p\in\mathcal{P}(\bm{x})}P_{p}(\bm{x})$ be a strongest path at $\bm{x}$. We define the main-path RM (MP-RM) as
@@
Let $\mathcal{P}*k(\bm{x})\subseteq\mathcal{P}(\bm{x})$ collect all paths whose interaction order does not exceed $k$, and assume $p^{\star}(\bm{x})\in\mathcal{P}*k(\bm{x})$. The multipath RM (MU-RM) is defined by aggregating all such paths:
@@
$\Delta U(\bm{x})\triangleq\sum_{p\in\mathcal{P}_{k}(\bm{x})\setminus{p^{\star}(\bm{x})}}P_{p}(\bm{x}).$ (20)
Since every summand in (20) is nonnegative, the MP-RM to MU-RM transition is pointwise monotone:
@@
and $\Delta U(\bm{x})=0$ holds if and only if no additional path contributes power at $\bm{x}$ within the adopted truncation $\mathcal{P}_{k}(\bm{x})$.
To describe spatial regularity, introduce a visibility indicator $V_{p}(\bm{x})\in{0,1}$ for each path $p$ and write
$P_{p}(\bm{x})=V_{p}(\bm{x}),\widetilde{P}_{p}(\bm{x}),$ (23)
@@
where $\widetilde{P}_{p}(\bm{x})$ captures the interaction losses and varies smoothly with $\bm{x}$ whenever the interaction sequence defining $p$ remains valid. Define the visibility state $S(\bm{x})={V_{p}(\bm{x})}_{p\in\mathcal{P}_{k}(\bm{x})}$ and partition $\Omega$ into regions of constant visibility state
$\mathcal{R}_{s}={\bm{x}\in\Omega:\ S(\bm{x})=s}.$ (25)
On each region $\mathcal{R}_{s}$, every $V_{p}(\bm{x})$ is constant, hence both $U_{\mathrm{MP}}$ and $U_{\mathrm{MU}}$ are finite sums of smooth functions $\widetilde{P}_{p}(\bm{x})$ and inherit the same smoothness within $\mathcal{R}_{s}$. Non-smooth behavior of spatial derivatives can only occur on the union of visibility boundaries $\partial\mathcal{R}_{s}$, where at least one $V_{p}(\bm{x})$ toggles. Under standard geometric optics visibility rules, these boundaries form a finite union of curves and thus occupy zero area measure in $\Omega$. Therefore, the MU-RM differs from the MP-RM through a nonnegative residual $\Delta U$ that is smooth almost everywhere and can change sharply only across visibility boundaries.
The decomposition (20) implies an additive relation for any linear integral operator. Let $\mathcal{L}_{\sigma}$ be a low-pass operator with bandwidth parameter $\sigma$ represented by a nonnegative kernel $\kappa_{\sigma}$ of unit mass:
$\mathcal{L}_{\sigma}[U](\bm{x})=\int_{\Omega}\kappa_{\sigma}(\bm{x}-\bm{y}),U(\bm{y}),\mathrm{d}\bm{y},$ (26)
$\int_{\mathbb{R}^{2}}\kappa_{\sigma}(\bm{u}),\mathrm{d}\bm{u}=1,\quad\kappa_{\sigma}(\bm{u})\geq 0.$ (27)
@@
A quantitative control of the low-frequency increment follows from standard convolution inequalities. Using Young’s inequality,
$|\mathcal{L}_{\sigma}[\Delta U]|_{2}\leq|\kappa_{\sigma}|_{1},|\Delta U|_{2}=|\Delta U|_{2},$ (29)
@@
$|\mathcal{L}_{\sigma}[\Delta U]|_{\infty}\leq|\kappa_{\sigma}|_{\infty},|\Delta U|_{1}.$ (30)
Suppose $\Delta U$ is effectively supported on a subset $\mathcal{S}\subseteq\Omega$ with small area and is bounded by $\Delta U(\bm{x})\leq U_{\max}$. Then
$|\Delta U|_{1}=\int_{\Omega}\Delta U(\bm{x}),\mathrm{d}\bm{x}=\int_{\mathcal{S}}\Delta U(\bm{x}),\mathrm{d}\bm{x}\leq U_{\max}|\mathcal{S}|,$ (31)
which combined with (30) yields
$|\mathcal{L}_{\sigma}[\Delta U]|_{\infty}\leq|\kappa_{\sigma}|_{\infty},U_{\max}|\mathcal{S}|.$ (32)
For typical low-pass kernels, $|\kappa_{\sigma}|_{\infty}$ decreases as $\sigma$ increases, so the coarse-scale effect of a spatially localized residual is attenuated by smoothing. This formalizes a central structural implication of (20): the MP-RM dominates the large-scale skeleton of the RM, while the MU-RM introduces structured positive increments that are concentrated in specific geometric neighborhoods where additional paths become visible.

### 26 Mar 2026 v3

@@ Abstract
Radio maps (RMs) provide spatially continuous propagation characterizations essential for 6G network planning, but high-fidelity RM construction remains challenging. Rigorous electromagnetic solvers incur prohibitive computational latency, while data-driven models demand massive labeled datasets and generalize poorly from simplified simulations to complex multipath environments. This paper proposes RadioDiff-FS, a few-shot diffusion framework that adapts a pretrained main-path generator to multipath-rich target domains with only a small number of high-fidelity samples. The adaptation is grounded in a theoretical decomposition of the multipath RM into a dominant main-path component and a directionally sparse residual. This decomposition shows that the cross-domain shift corresponds to a bounded and geometrically structured feature translation rather than an arbitrary distribution change. A direction-consistency loss (DCL) is then introduced to constrain diffusion score updates along physically plausible propagation directions, thereby suppressing phase-inconsistent artifacts that arise in the low-data regime. Experiments show that RadioDiff-FS reduces NMSE by 59.5% on static RMs and by 74.0% on dynamic RMs relative to the vanilla diffusion baseline, achieving an SSIM of 0.9752 and a PSNR of 36.37 dB under severely limited supervision. Even in a one-shot setting with a single target-domain sample per scene, RadioDiff-FS outperforms all fully supervised baselines, confirming that the directional constraint provides an effective inductive bias under extreme data scarcity. Code is available at https://github.com/UNIC-Lab/RadioDiff-FS.
@@ I Introduction
Figure 1: Illustration of the few-shot learning for RM construction.
@@
    We establish a rigorous theoretical decomposition of the multipath RM into a dominant main-path component and a directionally sparse residual. This formulation provides a physical inductive bias and demonstrate that the cross-domain transition is characterized by learnable geometric shifts rather than chaotic distribution changes, and lays the theoretical foundation for efficient few-shot adaptation.
@@
    We design a Direction-Consistency Loss that functions as a lightweight manifold projector within the diffusion process. By constraining score function updates to physically plausible propagation directions, DCL penalizes gradient components associated with mode hallucinations, thereby ensuring that synthesized maps maintain global spatial coherence while accurately recovering fine-grained diffraction and reflection details.
@@ II-A RM Construction Methods
Initial studies, such as RadioUNet [10], demonstrated that convolutional encoder–decoder architectures can produce accurate two-dimensional (2D) pathloss maps from structured urban layouts, significantly reducing inference latency compared with ray tracing. Subsequent works like PMNet improved generalization and sample efficiency through advanced vision architectures. Although these discriminative models remain computationally efficient and interpretable, they typically require dense supervision and are difficult to adapt to out-of-distribution environments or scenarios with sparse sensor coverage. Moreover, most of these methods focus on planar pathloss estimation, ignoring vertical structure and multi-metric coupling [27, 23]. To overcome the limitations of pointwise regression under sparse measurements, conditional generative models have been explored. RME-GAN [35] and DeepREM [2] propose using conditional generative adversarial networks and completion autoencoders to reconstruct dense RMs from limited measurements. These models integrate global propagation patterns and local texture variations to enhance estimation performance. While they significantly outperform interpolation baselines, they still rely heavily on simulation data and do not incorporate electromagnetic consistency into the learning process. Graph neural networks (GNNs) offer another route by encoding structural priors derived from building layouts and sensor topology [3]. Recent methods formulate sparse measurements as nodes and use edge structures informed by connectivity or spatial correlation to facilitate map completion. RadioGAT [11] extends this by capturing spatial–spectral dependencies across multiple frequency bands. These models offer improved robustness to sampling irregularities and can generalize under semi-supervised settings. However, without explicit electromagnetic constraints, their behavior around complex propagation features such as diffraction edges and shadow zones remains difficult to guarantee.
@@
Diffusion models have recently emerged as a powerful alternative due to their generative structure and sample flexibility. RadioDiff [23] models RM generation as a conditional diffusion process and introduces attention mechanisms and fast Fourier modules to enhance feature extraction in dynamic environments. RadioDiff-3D builds upon this by incorporating three-dimensional (3D) convolutional networks and utilizing the UrbanRadio3D dataset, which contains multi-height, multi-parameter RMs including pathloss, direction-of-arrival (DoA), and time-of-arrival (ToA) [24]. These models achieve state-of-the-art performance in high-dimensional map estimation. However, without embedding physical priors, they may fail to reconstruct physically meaningful structures in unseen domains, especially in low-data regimes. To address this, physics-informed models have been introduced to embed electromagnetic theory into neural estimators. PEFNet integrates the volume integral equation into the learning objective, constraining the relationship between incident and total electric fields. More advanced methods, such as PhyRMDM [8] and RadioDiff-$k^{2}$ [25], incorporate Helmholtz residuals, source terms, and boundary conditions into diffusion model training. RadioDiff-$k^{2}$ in particular establishes a theoretical correspondence between wave propagation singularities and negative wave-number domains, enabling the network to identify and reconstruct diffraction-sensitive structures via a two-stage generative process. These approaches significantly improve physical plausibility and generalization, but they often assume access to accurate simulations or well-calibrated measurements, which remain costly and difficult to obtain at scale. Task-specific strategies such as RM inpainting based on propagation priority and radio depth maps have also been proposed to address areas where sensing is physically restricted [20]. In parallel, the development of large-scale datasets such as UrbanRadio3D has greatly expanded the coverage of height, geometry, and multi-metric labeling, providing an essential benchmark for high-fidelity 3D×3D RM construction. Nevertheless, the cost and time required for data generation via ray tracing or controlled field measurements remain prohibitive in most deployment scenarios, and the resulting models often exhibit limited cross-region transferability.
@@ II-C Denoising Diffusion Model
where $t$ is the diffusion time index, and $f(\bm{x},t)$ denotes the drift term, $g(t)$ is a time-dependent diffusion coefficient, and $d\bm{w}$ represents a standard Wiener process. As $t$ increases, the perturbed distribution $p_{t}(\bm{x})$ asymptotically approaches an isotropic Gaussian. The generative process involves solving the corresponding reverse-time SDE:
@@
$s_{\theta}(\bm{x},t)=-\frac{\epsilon_{\theta}(\bm{x},t)}{\sqrt{1-\bar{\alpha}_{t}}}.$ (6)
Thus, DDPM can be viewed as a discretized instance of a score-based continuous model under a variance-preserving schedule.
@@ III System Model and Problem Formulation
We consider a wireless propagation scenario over a discretized planar region represented as an $N\times N$ uniform grid. Each cell in this grid corresponds to a spatial resolution element of width $\Delta$. The cell size is assumed small enough that the pathloss within it can be treated as constant. A RM over this region is defined as a matrix $\bm{P}\in\mathbb{R}^{N\times N}$, where each entry $P(i,j)$ denotes the pathloss at the center of cell $(i,j)$. The pathloss is measured at a fixed carrier frequency $f_{c}$ with the corresponding wavelength $\lambda=c/f_{c}$, where $c$ denotes the speed of light in free space. A single transmitter, modeled as a point radiator, is positioned at $\bm{r}=\langle d_{x},d_{y},d_{z}\rangle$, where $(d_{x},d_{y})$ denotes the horizontal coordinates in the 2D region and $d_{z}>0$ indicates the height above the ground. The transmitter emits electromagnetic energy isotropically or according to a specified antenna gain pattern, and serves as the sole radiation source. Although we focus on the single-transmitter setting, the formulation can be extended to multi-antenna or multi-site deployments by considering multiple independent transmitter locations.
The surrounding environment is composed entirely of static structures such as buildings. The geometric layout of these structures is captured by a binary occupancy matrix $\bm{H}\in\{0,1\}^{N\times N}$. An entry $H(i,j)=1$ indicates that the cell at location $(i,j)$ is fully occupied by a non-penetrable obstacle, while $H(i,j)=0$ represents free space. These obstacles are assumed to fully block direct wave propagation. Given a transmitter position $\bm{r}$ and a static environment $\bm{H}$, the goal is to estimate the pathloss field $\bm{P}_{\star}(\bm{r})=\mathcal{F}(\bm{H};\bm{r})$, where $\mathcal{F}$ denotes the environment-to-field operator. In this work, we consider two distinct types of RMs. The first type is the static RM (SRM), which models only the dominant propagation path and captures large-scale attenuation effects such as free-space spreading and shadowing due to occlusions. The second type is the multi-path RM (MRM), which provides a more detailed characterization of the propagation field by incorporating multipath effects. These effects include reflections and diffractions induced by the surrounding static geometry. The MRM offers a more physically complete description of the electromagnetic field, particularly in non-line-of-sight regions.
@@
To support the construction of these RMs, we assume the availability of two labeled datasets. The first dataset is denoted as $\mathcal{D}_{\mathrm{SRM}}=\{\bm{H}_{k},\bm{r}_{k},\bm{P}_{k}^{\mathrm{SRM}}\}_{k=1}^{K}$ and contains a large number of examples generated under the SRM setting. These samples are efficient to compute and cover a wide variety of layouts and transmitter placements. The second dataset is denoted as $\mathcal{D}_{\mathrm{MRM}}=\{\bm{H}_{\ell},\bm{r}_{\ell},\bm{P}_{\ell}^{\mathrm{MRM}}\}_{\ell=1}^{L}$ and contains a small number of high-fidelity samples that capture multipath propagation effects. Due to the computational overhead of accurate multipath modeling, the size of this dataset is limited, i.e., $L\ll K$. The objective of the above problem can be formulated as follows.
@@
$\min_{\bm{\theta}}\sum_{\ell=1}^{L}\frac{1}{L}\mathcal{L}\left(\bm{P}_{\ell}^{\mathrm{MRM}},\hat{\bm{P}}_{\ell}^{\mathrm{MRM}}\right),$ (15)
$\text{s.t.}\quad\hat{\bm{P}}_{\ell}^{\mathrm{MRM}}=\bm{\mu}_{\bm{\theta}}\left(\bm{H},\bm{r}\mid\mathcal{D}_{\mathrm{SRM}},\mathcal{D}_{\mathrm{MRM}}\right),$ (15a)
where $\bm{\theta}$ denotes the trainable parameters of the neural network $\bm{\mu}(\cdot)$.
@@ IV-A Main-Path Decomposition
Consider a bounded planar domain $\Omega\subset\mathbb{R}^{2}$ and a fixed transmitter–environment configuration. At a receiver location $\bm{x}\in\Omega$, let $\mathcal{P}(\bm{x})$ denote a finite set of propagation paths induced by the environment. Under the power-incoherent superposition model commonly adopted in RM construction, the received power field is expressed as the nonnegative sum
@@
where $P_{p}(\bm{x})$ aggregates geometric spreading and interaction-dependent losses along path pp. Let $p^{\star}(\bm{x})\in\arg\max_{p\in\mathcal{P}(\bm{x})}P_{p}(\bm{x})$ denote a strongest path at $\bm{x}$. The main-path RM (MP-RM) is then defined as
@@
Let $\mathcal{P}_{k}(\bm{x})\subseteq\mathcal{P}(\bm{x})$ denote the subset that collects all paths whose interaction order does not exceed $k$, and assume $p^{\star}(\bm{x})\in\mathcal{P}_{k}(\bm{x})$. The multipath RM (MU-RM) is defined by aggregating all such paths:
@@
$\Delta U(\bm{x})\triangleq\sum_{p\in\mathcal{P}_{k}(\bm{x})\setminus\{p^{\star}(\bm{x})\}}P_{p}(\bm{x}).$ (20)
Since every summand in (20) is nonnegative, the transition from MP-RM to MU-RM is pointwise monotone:
@@
and $\Delta U(\bm{x})=0$ holds if and only if no additional path contributes power at $\bm{x}$ within $\mathcal{P}_{k}(\bm{x})$.
To describe the spatial regularity of the residual, we introduce a visibility indicator $V_{p}(\bm{x})\in\{0,1\}$ for each path $p$ and factorize the per-path power as
$P_{p}(\bm{x})=V_{p}(\bm{x})\,\widetilde{P}_{p}(\bm{x}),$ (23)
@@
where $\widetilde{P}_{p}(\bm{x})$ captures the interaction losses along path pp. This factor varies smoothly with $\bm{x}$ whenever the interaction sequence defining $p$ remains valid. We define the visibility state as $S(\bm{x})=\{V_{p}(\bm{x})\}_{p\in\mathcal{P}_{k}(\bm{x})}$ and partition $\Omega$ into regions of constant visibility state
$\mathcal{R}_{s}=\{\bm{x}\in\Omega:\ S(\bm{x})=s\}.$ (25)
On each region $\mathcal{R}_{s}$, every indicator $V_{p}(\bm{x})$ is constant. Therefore, both $U_{\mathrm{MP}}$ and $U_{\mathrm{MU}}$ reduce to finite sums of smooth functions $\widetilde{P}_{p}(\bm{x})$ and inherit the same smoothness within $\mathcal{R}_{s}$. Non-smooth behavior of spatial derivatives can only occur on the union of visibility boundaries $\partial\mathcal{R}_{s}$, where at least one indicator $V_{p}(\bm{x})$ toggles. Under standard geometric optics rules, these boundaries form a finite union of curves and occupy zero area measure in $\Omega$. The MU-RM therefore differs from the MP-RM through a nonnegative residual $\Delta U$ that is smooth almost everywhere and changes sharply only across visibility boundaries. The decomposition (19) implies an additive relation under any linear integral operator. Let $\mathcal{L}_{\sigma}$ be a low-pass operator with bandwidth parameter $\sigma$, represented by a nonnegative kernel $\kappa_{\sigma}$ of unit mass:
$\mathcal{L}_{\sigma}[U](\bm{x})=\int_{\Omega}\kappa_{\sigma}(\bm{x}-\bm{y})\,U(\bm{y})\,\mathrm{d}\bm{y},$ (26)
$\int_{\mathbb{R}^{2}}\kappa_{\sigma}(\bm{u})\,\mathrm{d}\bm{u}=1,\quad\kappa_{\sigma}(\bm{u})\geq 0.$ (27)
@@
A quantitative bound on the low-frequency increment follows from standard convolution inequalities. We denote the $L^{p}$ norm of a function $f$ over $\Omega$ by $\|f\|_{p}$. Applying Young’s convolution inequality yields
$\|\mathcal{L}_{\sigma}[\Delta U]\|_{2}\leq\|\kappa_{\sigma}\|_{1}\,\|\Delta U\|_{2}=\|\Delta U\|_{2},$ (29)
@@
$\|\mathcal{L}_{\sigma}[\Delta U]\|_{\infty}\leq\|\kappa_{\sigma}\|_{\infty}\,\|\Delta U\|_{1}.$ (30)
Suppose $\Delta U$ is effectively supported on a subset $\mathcal{S}\subseteq\Omega$ with small area and is bounded by $\Delta U(\bm{x})\leq U_{\max}$. Then the $L^{1}$ norm of the residual satisfies
$\|\Delta U\|_{1}=\int_{\Omega}\Delta U(\bm{x})\,\mathrm{d}\bm{x}=\int_{\mathcal{S}}\Delta U(\bm{x})\,\mathrm{d}\bm{x}\leq U_{\max}\,|\mathcal{S}|,$ (31)
where $|\mathcal{S}|$ denotes the area of $\mathcal{S}$. Combining (31) with (30) yields
$\|\mathcal{L}_{\sigma}[\Delta U]\|_{\infty}\leq\|\kappa_{\sigma}\|_{\infty}\,U_{\max}\,|\mathcal{S}|.$ (32)
For typical low-pass kernels, $\|\kappa_{\sigma}\|_{\infty}$ decreases as $\sigma$ increases. Thus the coarse-scale effect of a spatially localized residual is attenuated by smoothing. This result formalizes a central structural implication of (19): the MP-RM dominates the large-scale skeleton of the RM, while the MU-RM introduces structured positive increments concentrated in geometric neighborhoods where additional paths become visible.